Related to Solved Problem #4] Suppose that real GDP is currently $13.1 trillion and potential real GDP is $14.0 trillion, or a gap of $900 billion. The government purchases multiplier is 5.0, and the tax multiplier is 4.0 Holding other factors constant, by how much will government purchases need to be increased to bring the economy to equilibrium at potential GDP? Government spending will need to be increased by Sllon. (Enter your response rounded to the nearest whole number.)
Answer #1

The government spending need to be increased by = (Potential real GDP - Real GDP) / Purchase multiplier

= (14-13.1)/5

= 900/5 = $180 billion